# Robust Multimodal Learning with Missing Modalities via Parameter-Efficient Adaptation

### v3

@@ 4.2 Implementation Details
To investigate missing modality adaptation performance in multimodal semantic and material segmentation tasks, we use the CMNeXt [70] as the base model. We use multimodal transformer [53] as the base model for multimodal sentiment analysis. We train the corresponding base model with all the input modalities for each dataset. To evaluate performance with missing modalities, we provide the available modalities and set the missing modalities to zero. To perform model adaptation for any modality subset $\mathcal{S}\subset\mathcal{M}$, we freeze the pretrained weights and insert learnable SSF layers. Then we fine tune the learnable parameters for 100 epochs for multimodal segmentation tasks and until convergence for multimodal sentiment analysis tasks.
@@
For multimodal segmentation tasks, we set the initial learning rate to $6\times 10^{-5}$ and applied a polynomial learning rate scheduler with a power of 0.9. The first 10 epochs were set as the warm-up, during which the learning rate was set to 0.1 times the original rate. The scale parameters $(\gamma)$ were initialized with all 1s and the shift parameters $(\beta)$ were initialized with all 0s. We used cross-entropy loss function and AdamW optimizer as proposed in [37], with an epsilon value of $10^{-8}$ and a weight decay of 0.01. We used a batch size of 4 and report single scale performance for all the datasets. All other hyper-parameters and configurations are the same as [70]. For multimodal sentiment analysis tasks, we used the default settings for the datasets as configured in the codebase [66]. We have included additional details for each dataset and experimental setup in the supplementary section.
4.3 Baseline Methods
We report experiments and results for different methods that are listed as follows. Pretrained model refers to the base model that is trained with all the available modalities. We provide zeros for the missing modalities. Modality Duplication means that one of the available modalities is used as a substitution for the missing modality. Dedicated training refers to independent models trained for each input modality combination. Adapted model refers to the model that is adapted using our approach for each input modality combination.
Different robust methods have been proposed for different multimodal tasks. We compare our method with the following methods: masking and recursive meshing based approach SpiderMesh [13], variational probabilistic fusion based approach VPFNet [33], modality discrepancy reduction approach MDRNet [72], knowledge distillation and robust training based approach CRM [48] for robust RGB-thermal semantic segmentation. Dynamic channel exchange based approach CEN [56], dynamic token selection/generation based approach TokenFusion [55], assymetric fusion based approach AsymFusion [57] for robust RGB-depth semantic segmentation.
@@ 4.3 Baseline Methods
For every task/dataset, we adopted the experimental setup used in the corresponding previous studies. We used the reported results from prior works where possible. It is important to note that, because of this criteria, some of the baseline methods may only be present in specific experiments depending on the availability of their reported numbers. We also perform detailed comparison of SSF with other parameter-efficient adaptation techniques. For all the experiments we follow the same setup suggested by the corresponding papers and report mean accuracy (mAcc), F1 score and mean intersection over union (mIoU) when available.
Ablation studies and comparison of different parameter-efficient adaptation methods show that SSF-based adaptation provides overall best performance. We summarize the results for scale only, shift only, BitFit [5], norm layer fine-tuning and LoRA [24] in Table 4. We also add detailed comparison for all the datasets in the supplementary section of the paper.
4.4 Experiments for Multimodal Segmentation
@@
In this section, we present experiments for multimodal semantic segmentation on MFNet and NYUDv2 datasets, and multimodal material segmentation on MCubeS dataset. First, we show an overall comparison of our approach with existing baselines on Table 1 and then we compare our approach with existing robust methods on Table 2 and 3. We report the detailed results for multimodal material segmentation along with per class % intersection over union (IoU) comparisons between the Pretrained and Adapted models in the supplementary section of the paper.
@@ 4.4.1 Overall performance comparison
We present the performance comparison of Pretrained, Modality Duplication, Dedicated, and Adapted networks for different missing modalities in Table 1. We observe that the performance of the Pretrained model drops significantly with missing modalities. We see a 6.39% drop when Thermal is missing on MFNet dataset and 5.11% drop when Depth is missing on NYUDv2 dataset compared to the case when all modalities are available. The effect is amplified when RGB gets missing as we observe 24.62% drop on MFNet dataset and 51.04% drop on NYUDv2 dataset. On MCubeS dataset, we observe 2.48–9.22% drop in pretrained model when different modality combinations are missing. Similar trend of performance drop is observed for modality duplication approach though it performs better than pretrained models for most of the cases.
@@
The overall performance of Adapted models with missing modalities is significantly better than Pretrained models and Modality Duplication approach. For MFNet, an improvement of 1.51% compared to the Pretrained model is observed when RGB is available and thermal is missing. The Adapted model performance is also close to the performance of Dedicated network trained for RGB only. The adapted model shows a significant improvement of 15.41% compared to the Pretrained model when RGB is missing. For NYUDv2 dataset, we see 1.63% and 31.46% performance improvement compared to Pretrained model when depth and RGB are missing, respectively. In both cases, the performance of the Adapted model is better than the Dedicated model. For all input combinations on MCubeS dataset, we see 1.82–8.11% performance improvement compared to the Pretrained model. The Adapted model performs better than Dedicated models in all the cases.
@@
Table 2: Performance comparison with existing robust methods for MFNet dataset. RGB and Thermal columns report performance when only RGB and Thermal are available. Average column reports average performance when one of the two modalities gets missing. ‘-’ indicates that results for those cells are not published. Mean accuracy (mAcc) and % mean intersection over union (mIoU) are shown for all the experiments.
Methods RGB Thermal Average
@@
FuseNet [19] 11.11 10.31 41.33 36.85 26.22 23.58
MFNet [17] 26.62 24.78 19.65 16.64 23.14 20.71
RTFNet [51] 44.89 37.30 26.41 24.57 35.65 30.94
SAGate [6] 32.01 30.57 13.34 12.51 22.68 21.54
FEANet [8] 15.96 8.69 58.35 48.72 37.16 28.71
MDRNet [72] 57.11 45.89 41.98 30.19 49.55 38.04
VPFNet [33] 48.14 41.08 42.20 35.80 45.17 38.44
SpiderMesh [13] - 39.60 - 50.50 - 45.05
CRM [48] - 52.70 - 53.10 - 52.90
Adapted (Ours) 67.18 55.22 66.70 50.89 66.94 53.06
Table 3: Performance comparison with existing robust methods for NYUDv2 dataset. RGB and Depth columns report performance when only RGB and Depth are available. Average column indicates average performance when one of the two modalities gets missing. $\ast$ indicates that available codes and pretrained models from the authors were used to generate the results. Other results are from the corresponding papers.
Methods RGB Depth Average
@@
FCN [36] 44.70 31.60 35.70 25.20 40.20 28.40
Dilated FCN-2s [27] 47.10 32.30 39.30 26.80 43.20 29.55
AsymFusion (R-101) [57] 59.00 46.50 45.60 34.30 52.30 40.40
CEN (R-101) [56] $\ast$ 51.77 39.59 28.98 19.32 40.38 29.46
TokenFusion [55] $\ast$ 63.49 49.32 46.83 36.84 55.16 43.08
Adapted (Ours) 67.96 52.82 52.42 36.72 60.19 44.77
4.4.2 Performance comparison for RGB-Thermal semantic segmentation on MFNet dataset
We compare the performance of the Adapted model with existing robust models for RGB-thermal semantic segmentation on MFNet dataset in Table 2. The results show that the Adapted model offers the best average performance compared to the other baseline models, in terms of mean accuracy and % mIoU. Among the robust models, robust training and knowledge distillation based model CRM [48] shows competitive performance with the Adapted model. The Adapted model performs better when only RGB is available while CRM performs better when only Thermal is available. Notably CRM is designed specifically for RGB-Thermal pairs and requires specialized training approach and utilizes self-distillation loss between the clean and masked modalities to train the model. In contrast, our approach is applicable to any input modalities and does not require any specialized training technique. Our approach performs significantly better compared to masking and recursive meshing based SpiderMesh [13], variational pobabilistic fusion based VPFNet [33] and modality discrepancy reduction based MDRNet [72] models.
4.4.3 Performance comparison for RGB-Depth semantic segmentation on NYUDv2 dataset
Table 3 shows the performance comparison with existing robust models for RGB-Depth semantic segmentation on NYUDv2 dataset. For our experiments, we use HHA-encoded images proposed by [16] instead of raw depth maps. On an average, the Adapted model performs better than the existing robust models. Dynamic token selection and substitution based model TokenFusion [55] performs slightly better (+0.12%) in terms of mIoU when Depth is available and RGB is missing, but shows larger drop (-5.59%) in mean accuracy. On the other hand, the Adapted model performs significantly better (+3.5% mIoU and +4.47% mean accuracy) when RGB is avalable and Depth is missing. The average performance of the Adapted model is also better than the TokenFusion model despite the fact that TokenFusion was designed to work with RGB-Depth pair, whereas the Adapted method is independent of input modalities. Our method also performs significantly better compared to dynamic channel exchange based CEN [56] and asymmetric fusion based AsymFusion [57] models.
@@ 4.4.4 Visualization of predictions
For qualitative analysis, we show some examples of the predicted segmentation maps form the Pretrained and Adapted models in Figure 2. For each dataset, we show the input images, predictions when all the modalities are available during both training and test time (CMNeXt column), predictions from the pretrained and adapted models for different available/missing modality scenarios (Available input modality names are shown in parentheses above each image). We see in Figure LABEL:fig:vis-mfnet-1, the Pretrained model fails to detect humans when only RGB images are available and cars when only Thermal images are available. The adapted model can detect both humans and cars with missing modalities.
For NYUDv2 dataset, as shown in Figure LABEL:fig:vis-nyu-1, the Adapted model can detect window, bed, and furniture with higher accuracy than the Pretrained model with missing modalities. We only show the RGB input images for MCubeS dataset for brevity in Figure LABEL:fig:vis-mcubes-1. The Adapted model can identify sand, sky, and gravel with higher accuracy than the pretrained model. In all cases, the predictions from the Adapted model with missing modalities are closer to the predictions of the pretrained model with all the input modalities. We provide additional visualizations in the supplementary section of the paper.
Table 4: Performance comparison (mIoU) of different parameter-efficient adaptation techniques for MFNet, NYUDv2, and MCubeS datasets. Each column reports mIoU of the Adapted model with the corresponding modalities, and Avg indicates average performance. A and D denote Angle and Degree of Linear Polarization.
Datasets MFNet NYUDv2 MCubeS
Methods RGB Thermal Avg RGB Depth Avg RGB RGB-A RGB-A-D Avg
Pretrained 53.71 35.48 44.60 51.19 5.26 28.23 42.32 48.81 49.06 46.73
Dedicated 55.86 53.34 54.60 52.18 33.49 42.84 48.16 48.42 49.48 48.69
Scale Only 54.77 49.23 52.00 53.04 36.12 44.58 50.16 50.55 51.13 50.61
Shift Only 54.57 48.96 51.77 53.04 36.25 44.65 50.13 50.40 50.86 50.46
BitFit 54.39 49.07 51.73 53.09 36.64 44.87 50.19 50.57 51.07 50.61
LoRA 54.19 47.45 50.82 52.87 34.97 43.92 49.59 50.07 50.80 50.15
Norm 54.65 47.49 51.07 53.05 34.73 43.49 49.95 50.51 51.07 50.51
Table 5: Comparison of our adaptation technique with existing methods for multimodal sentiment analysis.
Datasets Methods Audio Visual Audio-Visual Average
@@
CMU -MOSI MulT ([53]) 48.31 40.98 52.44 51.77 48.93 41.95 49.89 44.90
MFN ([68]) 56.86 44.81 55.95 42.94 56.86 51.07 56.56 46.27
TFN ([67]) 42.23 25.07 42.38 25.40 42.23 25.07 42.28 25.18
BERT_MAG ([43]) 57.77 42.31 57.77 42.31 57.77 42.31 57.77 42.31
LMF ([35]) 42.23 25.07 43.14 27.54 43.29 27.61 42.89 26.74
Adapted (Ours) 50.00 46.71 54.88 54.39 55.49 53.96 53.46 51.69
CMU - MOSEI MulT ([53]) 37.15 20.12 38.28 23.70 41.91 32.78 39.11 25.53
MFN ([68]) 58.48 58.31 60.35 59.48 59.74 60.37 59.52 59.39
TFN ([67]) 37.15 20.12 37.15 20.12 37.15 20.12 37.15 20.12
BERT_MAG ([43]) 62.83 48.50 61.39 49.70 62.83 48.51 62.35 48.90
LMF ([35]) 42.38 34.48 57.15 57.85 55.94 56.63 51.82 49.65
Adapted (Ours) 62.85 55.55 62.49 60.00 63.32 60.69 62.89 58.75
4.4.5 Performance comparison with parameter-efficient model adaption techniques
Apart from robust models, we also compare different parameter-efficient adaptation techniques. We summarize the results in Table 4. For MFNet dataset, SSF outperforms all the methods and performance is significantly better than the Pretrained model and close to the Dedicated models. For NYUDv2 and MCubeS datasets, the Adapted model performs better than both Pretrained and Dedicated models. These experiments also show that SSF performs better than other methods for most of the input modality combinations for all the datasets. We show a detailed comparison for each dataset in terms of mean accuracy, F1 score and % mIoU in the supplementary section of this paper.
4.5 Experiments for Multimodal Sentiment Analysis

### v4

@@ 4.2 Implementation Details
We use CMNeXt [61] as the base model for multimodal segmentation tasks, multimodal transformer [62] for multimodal sentiment analysis, UMDR [63] for multimodal action recognition and ViLT [64] for multimodal classification. We train the corresponding base model with all the input modalities for each dataset. To evaluate performance with missing modalities, we provide the available modalities and set the missing modalities to zero for images and empty string for texts. To perform model adaptation for any modality subset $\mathcal{S}\subset\mathcal{M}$, we fine tune the learnable parameters until convergence for all the tasks.
@@
For multimodal segmentation tasks, we set the learning rate to $6\times 10^{-5}$ and apply polynomial learning rate scheduler with power = 0.9. The first 10 epochs are warm-up epochs and the learning rate is set to 0.1 times the original rate. The scale $(\gamma)$ and shift $(\beta)$parameters were initialized with all 1s and 0s respectively. We use cross-entropy loss and AdamW optimizer [65], with $\epsilon=10^{-8}$ and weight decay = 0.01. We set batch size to 4 and report single scale performance. All other hyper-parameters are the same as [61]. For multimodal sentiment analysis, action recognition and classification tasks, we used the default settings from [66], [63] and [27] respectively. Please refer to Section S2 in the supplementary materials for additional details.
@@
For every task/dataset, we show the reported results from prior works where possible. It is important to note that, because of this criteria, some of the baseline methods may only be present in specific experiments depending on the availability of their reported numbers. We also perform detailed comparison of SSF with other parameter-efficient adaptation techniques which we discuss in Section S4 in the supplementary materials.
Table 1: Performance comparison with different baseline methods for multimodal semantic segmentation on MFNet and NYUDv2 datasets and multimodal material segmentation on MCubeS dataset. We use CMNeXt as the base model. Bold letters represent best results.
Dataset Input Missing Pretrained Modality Duplication Dedicated Adapted (Ours)
MFNet RGB-Thermal - 60.10 60.10 60.10 60.10
RGB Thermal 53.71 52.33 55.86 55.22
Thermal RGB 35.48 44.43 53.34 50.89
NYUDv2 RGB-Depth - 56.30 56.30 56.30 56.30
RGB Depth 51.19 46.19 52.18 52.82
Depth RGB 5.26 13.94 33.49 36.72
MCubeS RGB-AoLP-DoLP-NIR - 51.54 51.54 51.54 51.54
RGB-AoLP-DoLP NIR 49.06 49.93 49.48 51.11
RGB-AoLP DoLP-NIR 48.81 49.23 48.39 50.62
RGB AoLP-DoLP-NIR 42.32 48.96 48.11 50.43
4.3 Experiments on Multimodal Segmentation
@@
In this section, we present experimental results for multimodal semantic and material segmentation. First, we show an overall comparison of our approach with baselines methods and then we compare with existing robust methods.
@@ 4.3.1 Overall performance comparison
We report experimental results for different baseline methods in Table 1. Pretrained model refers to the base CMNeXt model trained with all the available modalities. Modality Duplication means that one of the available modalities is used as a substitution for the missing modality. Dedicated training indicates that we train one CMNeXt model for each input modality combination and use the model corresponding to the available modalities when some modalities get missing. Adapted model refers to the model that is adapted using our approach for each input modality combination.
@@
Pretrained model show significant performance drop with missing modalities. We see a 6.39% and 5.11% drop when Thermal is missing on MFNet and Depth is missing on NYUDv2, respectively, compared to the case when all modalities are available. The effect is amplified when RGB gets missing as we observe 24.62% and 51.04% drop on MFNet and NYUDv2 dataset respectively. On MCubeS dataset, we observe 2.48–9.22% drop in pretrained model when different modality combinations are missing. Similar trend of performance drop is observed for modality duplication approach though it performs better than pretrained models for most of the cases.
@@
The overall performance of the Adapted model is significantly better than Pretrained model and Modality Duplication approach. For MFNet, an improvement of 1.51% and 15.41% is observed compared to the Pretrained model when RGB and Thermal are available respectively. The performance of the Adapted models is also close to the Dedicated models. For NYUDv2 dataset, we see 1.63% and 31.46% performance improvement compared to Pretrained model when depth and RGB are missing, respectively. For all input combinations on MCubeS dataset, we see 1.82–8.11% performance improvement compared to the Pretrained model. The Adapted model performs better than Dedicated models on NYUDv2 and MCubeS datasets. Per-class IoU analysis shows that adapted models perform better than pretrained models for most of the classes which provides an overall performance improvement as discussed in Section S5.
Feature modulation during adaptation helps the model learn better feature representation and thus it performs better when modalities are missing. We discuss this in Section 4.7.1. Results also indicates that we do not need to train a dedicated network for each modality combination which requires more time and computation resources. Rather adapting one base model is sufficient to have comparable or even better performance in missing modality scenarios with less time and computational overhead.
4.3.2 Comparison with robust methods on MFNet dataset
We compare the performance of the Adapted model with existing robust models for RGB-thermal semantic segmentation on MFNet dataset in Table 2. Results show that the Adapted model offers the best average performance compared to existing baseline methods. Among the robust models, complementary random masking and knowledge distillation based model CRM [20] shows competitive performance with the Adapted model. The Adapted model performs better when only RGB is available while CRM performs better when only Thermal is available. Notably CRM is designed specifically for RGB-Thermal pairs and requires specialized training approach. In contrast, our approach is generic, applicable to any input modalities and does not require any specialized training technique. Our approach performs significantly better compared to partial masking and recursive meshing based SpiderMesh [32], variational pobabilistic fusion based VPFNet [35] and modality discrepancy reduction based MDRNet [67] models.
@@ 4.3.2 Comparison with robust methods on MFNet dataset
Table 2: Performance comparison with existing robust methods for MFNet dataset. RGB and Thermal columns report performance when only RGB and only Thermal are available. Average column reports average performance when one of the two modalities gets missing. ‘-’ indicates that results for those cells are not published. indicates that available code and pretrained models from the authors were used to generate the results.
Methods Backbone Parameters (M) RGB Thermal Average
@@
FuseNet [68] VGG-16 [69] - 11.11 10.31 41.33 36.85 26.22 23.58
MFNet [53] DCNN [70] 0.73 26.62 24.78 19.65 16.64 23.14 20.71
RTFNet [71] ResNet-152 [72] 254.51 44.89 37.30 26.41 24.57 35.65 30.94
SAGate [3] ResNet-50 [72] 110.85 32.01 30.57 13.34 12.51 22.68 21.54
FEANet [73] ResNet [72] - 15.96 8.69 58.35 48.72 37.16 28.71
MDRNet [67] ResNet-50 [72] 64.60 57.11 45.89 41.98 30.19 49.55 38.04
VPFNet [35] ResNet-50 [72] - 48.14 41.08 42.20 35.80 45.17 38.44
SpiderMesh [32] ResNet-152 [72] 151.81 - 39.60 - 50.50 - 45.05
CRM [20] Swin-S [74] 117.68 - 52.70 - 53.10 - 52.90
CMNeXt [61] MiT-B4 [75] 116.56 60.74 53.71 38.18 35.48 49.46 44.60
Adapted (Ours) MiT-B4 [75] 117.35 67.18 55.22 66.70 50.89 66.94 53.06
Table 3: Performance comparison with existing robust methods for NYUDv2 dataset. RGB and Depth columns report performance when only RGB and only Depth are available. Average column indicates average performance when one of the two modalities gets missing. indicates that available code and pretrained models from the authors were used to generate the results. Other results are from the corresponding papers.
Methods Backbone Parameters (M) RGB Depth Average
@@
FCN [76] VGG-16 [69] 134.00 44.70 31.60 35.70 25.20 40.20 28.40
Dilated FCN-2s [77] VGG-19 [69] 55.81 47.10 32.30 39.30 26.80 43.20 29.55
AsymFusion [78] ResNet-101 [72] 118.20 59.00 46.50 45.60 34.30 52.30 40.40
CEN [79] ResNet-101 [72] 118.20 51.77 39.59 28.98 19.32 40.38 29.46
TokenFusion [4] MiT-B3 [75] 45.92 63.49 49.32 46.83 36.84 55.16 43.08
CMNeXt [61] MiT-B4 [75] 116.56 64.10 51.19 8.30 5.26 36.20 28.23
Adapted (Ours) MiT-B4 [75] 117.35 67.96 52.82 52.42 36.72 60.19 44.77
Figure 2: Examples of predicted segmentation maps for the Pretrained and Adapted models. Title above each subimage shows method name (available modalities). CMNeXt column shows the predictions with all the modalities. Segmentation quality improves significantly after model adaptation for all input modality combinations. Green boxes highlight areas with salient differences in results (e.g., cars and humans missing in the Pretrained model with missing modalities but visible in the Adapted model). For MCubeS dataset, we only show RGB input images for brevity. A, D and N denote angle of linear polarization, degree of linear polarization, and near-infrared, respectively.
4.3.3 Comparison with robust methods on NYUDv2 dataset
Table 3 shows the performance comparison with existing robust models for RGB-Depth semantic segmentation on NYUDv2 dataset. On an average, the Adapted model performs better than the existing robust models. Dynamic token selection and substitution based model TokenFusion [4] performs slightly better (+0.12%) in mIoU when Depth is available and RGB is missing, but shows larger drop (-5.59%) in mean accuracy. On the other hand, the Adapted model performs significantly better (+3.5% mIoU and +4.47% mean accuracy) when RGB is available and Depth is missing. The average performance of the Adapted model is also better than the TokenFusion model despite the fact that TokenFusion was designed to work with RGB-Depth pair, whereas our approach is independent of input modalities. Our method also performs significantly better compared to dynamic channel exchange based CEN [79] and asymmetric fusion based AsymFusion [78] models.
@@ 4.3.4 Visualization of predictions
For qualitative analysis, we show some examples of the predicted segmentation maps form the Pretrained and Adapted models in Figure 2. For each dataset, we show the input images, predictions when all the modalities are available (CMNeXt column), predictions from the pretrained and adapted models for different available/missing modality scenarios (Available input modality names are shown in parentheses above each image). We see in Figure LABEL:fig:vis-mfnet-1, the Pretrained model fails to detect humans when only RGB images are available and cars when only Thermal images are available. The adapted model can detect both humans and cars with missing modalities.
On NYUDv2 dataset, as shown in Figure LABEL:fig:vis-nyu-1, the Adapted model can detect window, bed, and furniture with higher accuracy than the Pretrained model with missing modalities. On MCubeS dataset, the Adapted model can identify sand, sky, and gravel with higher accuracy than the pretrained model. In all cases, the predictions from the Adapted model with missing modalities are closer to the predictions of the pretrained model with all the input modalities. We provide additional visualizations in Figure S3 in the supplementary materials.
Table 4: Comparison of our adaptation technique with existing methods for multimodal sentiment analysis on CMU-MOSI and CMU-MOSEI datasets.
Datasets Methods Backbone Parameters (M) Audio Visual Audio-Visual Average
@@
CMU-MOSI MulT [62] Transformer [80] 2.58 48.31 40.98 52.44 51.77 48.93 41.95 49.89 44.90
MFN [81] LSTM [82] 2.17 56.86 44.81 55.95 42.94 56.86 51.07 56.56 46.27
TFN [83] LSTM [82] 5.04 42.23 25.07 42.38 25.40 42.23 25.07 42.28 25.18
BERT_MAG [84] BERT [85] 110.83 57.77 42.31 57.77 42.31 57.77 42.31 57.77 42.31
LMF [86] LSTM [82] 1.10 42.23 25.07 43.14 27.54 43.29 27.61 42.89 26.74
Adapted (Ours) Transformer [80] 2.60 50.00 46.71 54.88 54.39 55.49 53.96 53.46 51.69
CMU-MOSEI MulT [62] Transformer [80] 2.58 37.15 20.12 38.28 23.70 41.91 32.78 39.11 25.53
MFN [81] LSTM [82] 2.17 58.48 58.31 60.35 59.48 59.74 60.37 59.52 59.39
TFN [83] LSTM [82] 5.04 37.15 20.12 37.15 20.12 37.15 20.12 37.15 20.12
BERT_MAG [84] BERT [85] 110.83 62.83 48.50 61.39 49.70 62.83 48.51 62.35 48.90
LMF [86] LSTM [82] 1.10 42.38 34.48 57.15 57.85 55.94 56.63 51.82 49.65
Adapted (Ours) Transformer [80] 2.60 62.85 55.55 62.49 60.00 63.32 60.69 62.89 58.75
Table 5: Performance (top-1 accuracy) comparison with existing methods for action recognition on NTU RGB+D dataset. RGB and Depth columns report performance when only RGB and only Depth are available. Avg column indicates average performance. indicates that available code and pretrained models were used to generate the results.
Method Backbone RGB Depth Avg
Modality Distill. [87] ResNet-50 [72] 73.42 70.44 71.93
Luo et al. [88] ResNet-18 [72] + GRU [89] 89.50 87.50 88.50
DMCL [90] ResNet-18 [72] 83.61 80.56 82.09
Motion-RGBD [91] DSN + DTN [91] 90.30 92.70 91.50
ActionMAE [23] ResNet-34 [72] + Transformer [80] 84.50 90.50 87.50
UMDR [63] DSN + DTN [91] 90.47 93.99 92.23
Adapted (Ours) DSN + DTN [91] 91.53 94.29 92.91
4.4 Experiments on Multimodal Sentiment Analysis